# Gravitational wave polarization modes and the kinematical tensors in general relativity and beyond

### v1

@@ 2 Linearized kinematical tensors and polarization modes
$\theta=\theta^{\mathrm{M}}+\frac{1}{2}\dot{h}^{k}{}_{k}$, $\sigma=\sigma^{\mathrm{M}}+\frac{1}{2}\left(\dot{h}_{ij}-\frac{1}{3}\dot{h}^{k}{}_{k}\eta_{ij}\right)\mathrm{d}x^{i}\otimes\mathrm{d}x^{j},$ $\omega=\omega^{\mathrm{M}}+\frac{1}{2}\partial_{[i}h_{j]t}\mathrm{d}x^{i}\wedge\mathrm{d}x^{j},$ (24)
@@
$\sigma^{\mathrm{M}}=2\left(-v^{k}\partial_{(j}v_{k)}+\frac{1}{3}\partial_{k}v^{k}v_{j}\right)\mathrm{d}t\otimes_{\mathrm{s}}\mathrm{d}x^{j}+\left(\partial_{(i}v_{j)}-\frac{1}{3}\partial_{k}v^{k}\,\eta_{ij}\right)\mathrm{d}x^{i}\otimes\mathrm{d}x^{j},$ (25b)
@@
are purely kinematic contributions, in the sense that do not involve the gravitational field, present already in Minkowski spacetime. The residual terms in (24), do involve the metric perturbation, and contribute differently for gravitational waves of different metric theories, as will be shown in §IV.
@@
$K_{ab}=\frac{1}{2}\left(\ddot{h}_{ab}+\partial_{a}\partial_{b}h_{tt}\right)-\partial_{(a}\dot{h}_{b)t}.$ (26)
@@
where uppercase indices where used to denote spatial coordinates transverse to the spatial propagation direction of the wave, namely $(x^{I})=(x,y)$. Then, the only nonvanishing components of the polarization modes (13) yield simply
${K^{\mathrm{l}}}=K_{zz},$ ${K^{\mathrm{s}}}={K^{I}}_{I},$ ${K^{\mathrm{v}}}_{J}=K_{zJ},$ ${K^{\mathrm{tt}}}_{IJ}=K_{IJ}-\frac{{K^{L}}_{L}}{2}\eta_{IJ},$ (29)
which coincides with those reported in the literature [cf., e.g., equation (6) of [15]]. Explicitly in terms of the perturbation one obtains
@@
${K^{\mathrm{tt}}}_{IJ}=\frac{1}{2}\left(\ddot{h}_{IJ}-\frac{1}{2}\ddot{h}^{K}{}_{K}\eta_{IJ}\right)+\frac{1}{2}\left(\partial_{I}\partial_{J}h_{tt}-\frac{1}{2}\partial^{K}\partial_{K}h_{tt}\eta_{IJ}\right)-\left(\partial_{(I}\dot{h}_{J)t}-\frac{1}{2}\partial^{K}\dot{h}_{Kt}\eta_{IJ}\right).$ (30d)
@@ 1 General relativity
where $\Box_{\eta}=\partial^{c}\partial_{c}$. Moreover, the residual coordinate freedom within the de Donder gauge, allows one to impose, also without loss of generality, that (cf. §4.4b of [16] or §35.4 of [13])
$h=0,$ $h_{at}=0,$ (34)
which together with (32) constitute the transverse-traceless (TT) gauge, $h_{ab}=h^{\mathrm{TT}}_{ab}$. This perturbation corresponds to a massless spin-2 field (cf. §III of [17]). Note that, despite the similarity in the names and the labels of $h^{\mathrm{TT}}_{ab}$ and ${K^{\mathrm{tt}}}_{ab}$, they are related to different type of decompositions of symmetric tensors: the former involves differential relations (32), while the latter is defined in a purely algebraic manner (13). A thorough discussion on this matter can be found in [18].
@@
$e^{+}=\mathrm{d}x\otimes\mathrm{d}x-\mathrm{d}x\otimes\mathrm{d}y,$ $e^{\times}=2\mathrm{d}x\otimes_{\mathrm{s}}\mathrm{d}y,$ (36)
@@
$h_{+}=\real\left(A_{+}\mathrm{e}^{-\mathrm{i}\Omega(t-z)}\right),$ $h_{\times}=\real\left(A_{\times}\mathrm{e}^{-\mathrm{i}\Omega(t-z)}\right),$ (37)
@@ 2 Scalar curvature dependent Lagrangian
Next, we consider an $f(R)$ theory, namely one determined by a Lagrangian that is a function of the scalar curvature
@@
$\varphi=\real\left(\alpha\mathrm{e}^{-\mathrm{i}(\varpi t-qz)}\right),$ $\varpi^{2}-q^{2}=\mu^{2},$ (44)
@@
For this $f(R)$ gravitational wave, stationary particles in this “generalized” TT frame do not fall freely, since the right hand side of (23) does not vanish, in contrast with GR. The corresponding kinematical tensors give
@@
${K^{\mathrm{l}}}=-\mu^{2}\varphi,$ ${K^{\mathrm{s}}}=2\ddot{\varphi},$ ${K^{\mathrm{tt}}}_{ab}=\frac{1}{2}\ddot{\tilde{h}}^{\mathrm{TT}}_{IJ},$ (46)
and ${K^{\mathrm{v}}}_{a}=0$. In addition to the contribution from $\tilde{h}^{\mathrm{TT}}_{IJ}$, which certainly is the same as that of GR, one observes that $\varphi$ contributes to the expansion, and to the longitudinal and transverse scalar modes.
@@ 3 Einstein-Bach gravity
$S_{\gamma}[g_{ab}]=\frac{1}{16\pi G}\int\mathrm{d}^{4}x\sqrt{-g}\left(R-\gamma C_{abcd}C^{abcd}\right).$ (47)
The contribution of the Weyl squared term to the field equations is the Bach tensor [22], that is traceless. Thus, the trace of the vacuum field equations yield $R=0$ in general. This theory is in fact a particular class of the theories whose Lagrangians are at most quadratic in the Riemann tensor [23]. We will assume $\gamma>0$ for causal wave propagation [cf. (49) below].
@@
$\psi_{+}=\real\left(B_{+}\mathrm{e}^{-\mathrm{i}(\varpi t-qz)}\right),$ $\psi_{\times}=\real\left(B_{\times}\mathrm{e}^{-\mathrm{i}(\varpi t-qz)}\right),$ $\psi_{zj}=\real\left(B_{j}\mathrm{e}^{-\mathrm{i}(\varpi t-qz)}\right),$ $\varpi^{2}-q^{2}=m^{2},$ (50)
@@
$\psi_{tt}=\left(\frac{q}{\varpi}\right)^{2}\psi_{zz},$ $\psi_{tj}=-\frac{q}{\varpi}\psi_{zj},$ $\psi_{IJ}=\psi^{\mathrm{TT}}_{IJ}-\frac{1}{2}\left(\frac{m}{\varpi}\right)^{2}\psi_{zz}\eta_{IJ},$ $\psi^{\mathrm{TT}}_{IJ}=\psi_{+}e^{+}_{IJ}+\psi_{\times}e^{\times}_{IJ}.$ (51)
As with the $f(R)$ case, this gravitational wave does not allow stationary particles in this “generalized” TT frame to fall freely [cf. (23)]. The kinematical tensors give
$\theta=\theta^{\mathrm{M}}+\frac{1}{2}\left(\frac{q}{\varpi}\right)^{2}\dot{\psi}_{zz},$ (52)
$\sigma=\sigma^{\mathrm{M}}+\frac{1}{2}\left(\dot{\tilde{h}}^{\mathrm{TT}}_{IJ}+\dot{\psi}^{\mathrm{TT}}_{IJ}\right)\mathrm{d}x^{I}\otimes\mathrm{d}x^{J}+\frac{m^{2}+2\varpi^{2}}{12\varpi^{2}}\dot{\psi}_{zz}\left(-\eta_{IJ}\mathrm{d}x^{I}\otimes\mathrm{d}x^{J}+2\mathrm{d}z\otimes\mathrm{d}z\right)+\dot{\psi}_{zJ}\mathrm{d}z\otimes\mathrm{d}x^{J},$ (53)
$\omega=\omega^{\mathrm{M}}+\frac{1}{2}\left(\frac{q}{\varpi}\right)^{2}\dot{\psi}_{zJ}\mathrm{d}z\wedge\mathrm{d}x^{J},$ (54)
and the polarization modes yield
${K^{\mathrm{l}}}=-\frac{1}{2}\frac{m^{4}}{\varpi^{2}}\psi_{zz},$ ${K^{\mathrm{s}}}=\frac{m^{2}}{2}\psi_{zz},$ ${K^{\mathrm{v}}}_{J}=-\frac{m^{2}}{2}\psi_{zJ},$ ${K^{\mathrm{tt}}}_{IJ}=\frac{1}{2}\left(\ddot{\tilde{h}}^{\mathrm{TT}}_{IJ}+\ddot{\psi}^{\mathrm{TT}}_{IJ}\right).$ (55)
Additionally to the GR contribution from $\tilde{h}^{\mathrm{TT}}_{IJ}$, the massive spin-2 field $\psi_{ab}$ contributes to all the kinematical tensors and to all the polarization modes: its TT part has a contribution similar to that of GR; its longitudinal component contribute to the expansion, to the diagonal components of the shear, and to the longitudinal and transverse scalar modes; and its longitudinal-transverse components contribute to the longitudinal-transverse components of both the shear and the vorticity, and to the vector mode. It stands out that the contributions to the vorticity and to the vector mode are nonvanishing.
@@ V Discussion
We have considered a cloud of freely falling test particles acted upon a gravitational wave, and studied relations between the kinematical tensors of the cloud and the polarization modes of the wave, which, to our knowledge, had not been studied before. The polarization modes of a gravitational wave with a definite spacetime propagation direction were expressed in terms of the kinematical tensors of the cloud in §2. In addition to the foreseen relations, the obtained expressions revealed that: the expansion contributes to the longitudinal mode; while both the shear and the vorticity contribute to the transverse scalar mode via some of their transverse components, and to the vector mode through some of their longitudinal-transverse components. This analysis was general, and only assumed the existence of a wave propagation direction, having the advantage to be applied to a wide variety of gravitational wave phenomena. More concrete aspects of gravitational waves might be incorporated and would serve to obtain more particular and more relevant relations, which would be interesting to explore in the future.
@@
Additionally, we have calculated the kinematical tensors and the polarization modes, for a cloud of slowly moving test particles that fall freely in an arbitrary linearized field. Being interested in studying gravitational waves beyond GR, made it necessary to consider nonstationary particles, since not all theories allow the particles to be stationary in the laboratory frame, as GR does [cf. (23)]. Linearized gravitational waves of certain metric theories were shown in §IV to exhibit most of the anticipated interrelationships between the kinematical tensors and the polarization modes. The GR wave contributes to the transverse components of the shear and to the transverse tensor mode. The $f(R)$ wave contributes additionally to the expansion, and to the longitudinal and transverse scalar modes. Finally, the Einstein-Bach wave contributes additionally to all the kinematical tensors and all the polarization modes, standing out nonvanishing contributions to the vorticity and to the vector mode.
@@
It is reasonable to inquiry about the relevance of the considered metric theories of gravity. It turns out their linearized gravitational waves are representative of the larger class of metric theories determined by a Lagrangian that is a general function of the Riemann tensor. That this is so, comes from the fact that to obtain the linearized field equations, terms in the Lagrangian up to second degree in the Riemann tensor are the only ones that contribute (cf. §II of [20]), which can be chosen to be the quadratic term of an $f(R)$ and the Weyl squared scalar considered in §2 and §3, respectively. Moreover, the $f(R)$ example turns out to be representative of scalar-tensor theories, since their gravitational waves are essentially the same (cf. §3 of [27]).
@@
Our general expressions for the linearized kinematical tensors and the polarization modes were written in terms of the components of the metric perturbation, which facilitates their evaluation on linearized gravitational waves of metric theories of gravity found in the literature. Although it has the disadvantage of being written in terms of gauge-dependent variables, the gravitational waves of the examples we considered were expressed in a completely fixed gauge, thus excluding nonphysical coordinate effects. Rewriting our expressions in terms of gauge-independent metric potentials (cf., e.g., §2.2 of [28] or §III of [4]), specially for the kinematical tensors, would ease their interpretation for general weak gravitational waves, but this goes beyond the scope of this work.

### v2

@@ 2 Linearized kinematical tensors and polarization modes
$\theta=\theta^{\mathrm{M}}+\frac{1}{2}\dot{h}^{k}{}_{k}$, (23a)
$\sigma=\sigma^{\mathrm{M}}+\frac{1}{2}\left(\dot{h}_{ij}-\frac{1}{3}\dot{h}^{k}{}_{k}\eta_{ij}\right)\mathrm{d}x^{i}\otimes\mathrm{d}x^{j},$ (23b)
$\omega=\omega^{\mathrm{M}}+\frac{1}{2}\partial_{[i}h_{j]t}\mathrm{d}x^{i}\wedge\mathrm{d}x^{j},$ (23c)
@@
$\sigma^{\mathrm{M}}$ $=2\left(-v^{k}\partial_{(j}v_{k)}+\frac{1}{3}\partial_{k}v^{k}v_{j}\right)\mathrm{d}t\otimes_{\mathrm{s}}\mathrm{d}x^{j}$
$\quad+\left(\partial_{(i}v_{j)}-\frac{1}{3}\partial_{k}v^{k}\,\eta_{ij}\right)\mathrm{d}x^{i}\otimes\mathrm{d}x^{j},$ (24b)
@@
are purely kinematic contributions, in the sense that do not involve explicitly the gravitational field, present already in Minkowski spacetime. [^5] Despite the fact that these purely kinematical quantities do not exhibit explicitly the perturbation $h_{ab}$, they should not be regarded as independent of it, since the freely falling condition (22) in fact determines the velocities, and therefore those quantities, in terms of the perturbation. The residual terms in (23), do involve explicitly the metric perturbation, and contribute differently for gravitational waves of different metric theories, as will be shown in §IV. It is worth recalling that, for a given cloud of test particles, the kinematical tensors are uniquely defined by covariant equations [cf. (4)–(5)] and are therefore well-defined independently of the specific coordinate system one uses; thus, $\theta$, $\sigma$, and $\omega$ are in particular gauge-invariant within the weak field approximation.
The reader might wonder if one could use a coordinate system in which the purely kinematical contributions (24) simplify. In fact, since the particles are falling freely, one can use a Fermi normal coordinate system [26, 24, 25] adapted to the worldline of one reference particle which thus will remain at rest with respect to this coordinate system. This would simplify slightly the expressions in (24). However, we prefer not to restrict the coordinate system this way and rather restrict it by gauge conditions that simplify the form of the perturbation for the gravitational wave, as will be shown in the examples of §IV.
@@
$K_{ab}=\frac{1}{2}\left(\ddot{h}_{ab}+\partial_{a}\partial_{b}h_{tt}\right)-\partial_{(a}\dot{h}_{b)t},$ (25)
from which it follows that
$K_{ab}={\delta^{i}}_{a}{\delta^{j}}_{b}K_{ij}.$ (26)
In fact, a lengthy but direct calculation reveals that substituting the expressions for the linearized kinematical tensors (23)–(24) into (10), while taking into account the freely falling condition (22), yields an expression for $K_{ab}$ where all the terms involving the velocities of the particles cancel out, obtaining exactly the same expression as in (25).
@@
where uppercase indices were used to denote spatial coordinates transverse to the spatial propagation direction of the wave, namely $(x^{I})=(x,y)$. Then, the only nonvanishing components of the polarization modes (13) yield simply
${K^{\mathrm{l}}}=K_{zz},$ (29a)
${K^{\mathrm{s}}}={K^{I}}_{I},$ (29b)
$K^{\mathrm{v}}_{J}=K_{zJ},$ (29c)
$K^{\mathrm{tt}}_{IJ}=K_{IJ}-\frac{{K^{L}}_{L}}{2}\eta_{IJ}.$ (29d)
These expressions coincide with those reported in the literature [cf., e.g., equation (6) of [27]]. Explicitly in terms of the perturbation one obtains
@@
$K^{\mathrm{tt}}_{IJ}$ $=\frac{1}{2}\left(\ddot{h}_{IJ}-\frac{1}{2}\ddot{h}^{K}{}_{K}\eta_{IJ}\right)+\frac{1}{2}\left(\partial_{I}\partial_{J}h_{tt}-\frac{1}{2}\partial^{K}\partial_{K}h_{tt}\eta_{IJ}\right)$
$\quad-\left(\partial_{(I}\dot{h}_{J)t}-\frac{1}{2}\partial^{K}\dot{h}_{Kt}\eta_{IJ}\right).$ (30d)
@@ 1 General relativity
where $\Box_{\eta}=\partial^{c}\partial_{c}$. Moreover, the residual coordinate freedom within the de Donder gauge allows one to impose, also without loss of generality, that (cf. §4.4b of [28] or §35.4 of [23])
$h=0,$ $h_{at}=0.$ (34)
Taken together, (32) and (34) define the transverse-traceless (TT) gauge, $h_{ab}=h^{\mathrm{TT}}_{ab}$. This perturbation corresponds to a massless spin-2 field (cf. §III of [29]). Note that, despite the similarity in the names and the labels of $h^{\mathrm{TT}}_{ab}$ and $K^{\mathrm{tt}}_{ab}$, they are related to different type of decompositions of symmetric tensors: the former involves differential relations (32), while the latter is defined in a purely algebraic manner (13). A thorough discussion on this matter can be found in [30].
@@
$e^{+}=\mathrm{d}x\otimes\mathrm{d}x-\mathrm{d}y\otimes\mathrm{d}y,$ $e^{\times}=2\mathrm{d}x\otimes_{\mathrm{s}}\mathrm{d}y,$ (36)
@@
$h_{+}=\left(A_{+}\mathrm{e}^{-\mathrm{i}\Omega(t-z)}\right),$ $h_{\times}=\left(A_{\times}\mathrm{e}^{-\mathrm{i}\Omega(t-z)}\right),$ (37)
@@ 2 Scalar curvature dependent Lagrangian
Next, we consider an $f(R)$ theory, namely one determined by a Lagrangian that is a function of the scalar curvature $R$
@@
$\varphi=\left(\alpha\mathrm{e}^{-\mathrm{i}(\varpi t-qz)}\right),$ $\varpi^{2}-q^{2}=\mu^{2},$ (44)
@@
For this $f(R)$ gravitational wave, stationary particles in this “generalized” TT frame do not fall freely, since the right hand side of (22) does not vanish, in contrast with GR. Then, for freely falling particles, the expressions in (23) lead to the following kinematical tensors
@@
${K^{\mathrm{l}}}=-\mu^{2}\varphi,$ ${K^{\mathrm{s}}}=2\ddot{\varphi},$ $K^{\mathrm{tt}}_{IJ}=\frac{1}{2}\ddot{\tilde{h}}^{\mathrm{TT}}_{IJ},$ (46)
and $K^{\mathrm{v}}_{a}=0$. In addition to the contribution from $\tilde{h}^{\mathrm{TT}}_{IJ}$, which certainly is the same as that of GR, one observes that $\varphi$ contributes to the expansion, and to the longitudinal and transverse scalar modes (cf. table 1).
Table 1: Fields in the metric perturbation corresponding to linearized waves of the metric theories considered in §IV and their contribution to both the kinematical tensors and the polarization modes are shown. The field $\tilde{h}^{\mathrm{TT}}_{ab}$ is the spin-2 massless field of GR, the field $\varphi$ is the spin-0 massive field arising in $f(R)$ theories (cf. §2), and the field $\psi_{ab}$ is the spin-2 massive field appearing in Einstein-Bach gravity (cf. §3).
Field Spin Mass Contributes to
$\tilde{h}^{\mathrm{TT}}_{ab}$ 2 =0=0 $\sigma_{IJ}$, $K^{\mathrm{tt}}_{IJ}$
$\varphi$ 0 $\neq 0$ $\theta$, $K^{\mathrm{l}}$}, $K^{\mathrm{s}}$}
$\psi^{\mathrm{TT}}_{IJ}$ $\sigma_{IJ}$, $K^{\mathrm{tt}}_{IJ}$
$\psi_{zz}$ 2 $\neq 0$ $\theta$, $K^{\mathrm{l}}$}, $K^{\mathrm{s}}$}, $\sigma_{ii}$
$\psi_{zJ}$ $\sigma_{zJ}$, $\omega_{zJ}$, $K^{\mathrm{v}}_{J}$
@@ 3 Einstein-Bach gravity
$S_{\gamma}[g_{ab}]=\frac{1}{16\pi G}\int\mathrm{d}^{4}x\sqrt{-g}\left(R-\gamma C_{abcd}C^{abcd}\right),$ (47)
with $C_{abcd}$ the Weyl tensor and $\gamma$ a constant. The contribution of the Weyl squared term to the field equations is the Bach tensor [34], which is traceless. Thus, the trace of the vacuum field equations yields $R=0$ in general. This theory is in fact a particular class of the theories whose Lagrangians are at most quadratic in the Riemann tensor [35]. We will assume $\gamma>0$ for causal wave propagation [cf. (49) below].
@@
$\psi_{+}=\left(B_{+}\mathrm{e}^{-\mathrm{i}(\varpi t-qz)}\right),$ (50a)
$\psi_{\times}=\left(B_{\times}\mathrm{e}^{-\mathrm{i}(\varpi t-qz)}\right),$ (50b)
$\psi_{zj}=\left(B_{j}\mathrm{e}^{-\mathrm{i}(\varpi t-qz)}\right),$ (50c)
$\varpi^{2}-q^{2}=m^{2},$ (50d)
@@
$\psi_{tt}=\left(\frac{q}{\varpi}\right)^{2}\psi_{zz},$ (51a)
$\psi_{tj}=-\frac{q}{\varpi}\psi_{zj},$ (51b)
$\psi_{IJ}=\psi^{\mathrm{TT}}_{IJ}-\frac{1}{2}\left(\frac{m}{\varpi}\right)^{2}\psi_{zz}\eta_{IJ},$ (51c)
$\psi^{\mathrm{TT}}_{IJ}=\psi_{+}e^{+}_{IJ}+\psi_{\times}e^{\times}_{IJ}.$ (51d)
As with the $f(R)$ case, this gravitational wave does not allow stationary particles in this “generalized” TT frame to fall freely [cf. (22)]. The kinematical tensors (23) give
$\theta$ $=\theta^{\mathrm{M}}+\frac{1}{2}\left(\frac{q}{\varpi}\right)^{2}\dot{\psi}_{zz},$ (52a)
$\sigma$ $=\sigma^{\mathrm{M}}+\frac{1}{2}\left(\dot{\tilde{h}}^{\mathrm{TT}}_{IJ}+\dot{\psi}^{\mathrm{TT}}_{IJ}\right)\mathrm{d}x^{I}\otimes\mathrm{d}x^{J}$
$\quad+\frac{m^{2}+2\varpi^{2}}{12\varpi^{2}}\dot{\psi}_{zz}\left(-\eta_{IJ}\mathrm{d}x^{I}\otimes\mathrm{d}x^{J}+2\mathrm{d}z\otimes\mathrm{d}z\right)$
$\quad+\dot{\psi}_{zJ}\mathrm{d}z\otimes_{\mathrm{s}}\mathrm{d}x^{J},$ (52b)
$\omega$ $=\omega^{\mathrm{M}}+\frac{1}{2}\left(\frac{q}{\varpi}\right)^{2}\dot{\psi}_{zJ}\mathrm{d}z\wedge\mathrm{d}x^{J},$ (52c)
where $\partial_{z}\psi_{ab}=-(q/\varpi)\dot{\psi}_{ab}$ was used, while the polarization modes yield
${K^{\mathrm{l}}}=-\frac{1}{2}\frac{m^{4}}{\varpi^{2}}\psi_{zz},$ (53a)
${K^{\mathrm{s}}}=\frac{m^{2}}{2}\psi_{zz},$ (53b)
$K^{\mathrm{v}}_{J}=-\frac{m^{2}}{2}\psi_{zJ},$ (53c)
$K^{\mathrm{tt}}_{IJ}=\frac{1}{2}\left(\ddot{\tilde{h}}^{\mathrm{TT}}_{IJ}+\ddot{\psi}^{\mathrm{TT}}_{IJ}\right).$ (53d)
In addition to the GR contribution from $\tilde{h}^{\mathrm{TT}}_{IJ}$, the massive spin-2 field $\psi_{ab}$ contributes to all the kinematical tensors and to all the polarization modes: its TT part has a contribution similar to that of GR; its longitudinal component contributes to the expansion, to the diagonal components of the shear, and to the longitudinal and transverse scalar modes; and its longitudinal-transverse components contribute to the longitudinal-transverse components of both the shear and the vorticity, and to the vector mode. It stands out that the contributions to the vorticity and to the vector mode are nonvanishing. These contributions are summarized in table 1.
@@ V Discussion
We have considered a cloud of freely falling test particles acted upon by a gravitational wave, and studied relations between the kinematical tensors of the cloud and the polarization modes of the wave, which, to our knowledge, had not been studied before. The polarization modes of a gravitational wave with a definite spacetime propagation direction were expressed in terms of the kinematical tensors of the cloud in §2. In addition to the foreseen relations (cf. §I), the obtained expressions revealed that: the expansion contributes to the longitudinal mode; while both the shear and the vorticity contribute to the transverse scalar mode via some of their transverse components, and to the vector mode through some of their longitudinal-transverse components. This analysis was general, and only assumed a definite wave propagation direction, having the advantage of being applicable to a wide variety of gravitational wave phenomena. More concrete aspects of gravitational waves might be incorporated, serving to obtain more specific and relevant relations, which would be interesting to explore in the future.
@@
Additionally, we have calculated the kinematical tensors and the polarization modes for a cloud of slowly moving test particles that fall freely in an arbitrary linearized field. Being interested in studying gravitational waves beyond GR made it necessary to consider nonstationary particles, since not all theories allow the particles to be stationary in the laboratory frame, as GR does [cf. (22)]. Linearized gravitational waves of certain metric theories were shown in §IV to exhibit most of the anticipated relations between the kinematical tensors and the polarization modes. The GR wave contributes to the transverse components of the shear and to the transverse tensor mode. The $f(R)$ wave contributes additionally to the expansion, and to the longitudinal and transverse scalar modes. Finally, the Einstein-Bach wave contributes additionally to all the kinematical tensors and to all the polarization modes, standing out the nonvanishing contributions to the vorticity and to the vector mode. The contributions of each of these waves is summarized in table 1.
The inclusion of the kinematical tensors to the analysis of gravitational waves put forward in this work, provides new insights into the interpretation of the polarization modes, and offers a fresh perspective on devising new ways to find deviations from GR, both at the theoretical and at the phenomenological level.
@@
It is reasonable to inquire about the relevance of the considered metric theories of gravity. It turns out that their linearized gravitational waves are representative of those of the larger class of metric theories determined by a Lagrangian that is a general function of the Riemann tensor. That this is so, comes from the fact that to obtain the linearized field equations, terms in the Lagrangian up to second degree in the Riemann tensor are the only ones that contribute (cf. §II of [32]), which can be chosen to be the quadratic term of an $f(R)$ and the Weyl squared scalar considered in §2 and §3, respectively. Moreover, the $f(R)$ example turns out to be representative of second-order scalar-tensor theories, since their gravitational waves are essentially the same (cf. §3 of [4]). Hence, these results are applicable to a wide variety of metric theories of gravity.
@@
Our general expressions for the linearized kinematical tensors and the polarization modes were written in terms of the components of the metric perturbation, which facilitates their evaluation on linearized gravitational waves of metric theories of gravity found in the literature. Although it has the disadvantage of being written in terms of gauge-dependent variables, the gravitational waves of the examples we considered were expressed in a completely fixed gauge, thus excluding nonphysical coordinate effects. Rewriting our expressions in terms of gauge-independent metric potentials (cf., e.g., §2.2 of [39] or §III of [7]), especially for the kinematical tensors, would ease their interpretation for general weak gravitational waves, but this goes beyond the scope of this work.